an agronomist believes that a newly developed fertilizer will increase the mean harvest of eggplants by more than 2.5 kg. twenty six plants are treated with fertilizer and have a mean of 10.5 kg with standard deviation of 1.2. it is known that the population mean was 7.5 kg. test the claim at 0.01 level of significance  


1
Expert's answer
2022-05-04T01:41:32-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le10"

"H_a:\mu>10"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=25" degrees of freedom, and the critical value for a right-tailed test is "t_c = 2.485107."

The rejection region for this right-tailed test is "R = \{t: t > 2.485107\}"

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{10.5-10}{1.2/\sqrt{26}}\approx2.1246"

Since it is observed that "t = 2.1246 \le 2.485107=t_c ," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach: The p-value for right-tailed, "df=25" degrees of freedom, "t=2.1246" is "0.021843," and since "p= 0.021843>0.01=\alpha,"

it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is greater than 10, at the "\alpha = 0.01" significance level.

Therefore, there is not enough evidence to claim that a newly developed fertilizer will increase the mean harvest of eggplants by more than 2.5 kg, at the "\alpha = 0.01" significance level.